OpenStudy (anonymous):

how do you compute (d/dx) [cos^2(x) - cos(x^2)] so that it equals -2cos(x) sin(x) + 2xsin(x^2) when i do it i get zero

OpenStudy (anonymous):

What do you get when you chain rule the first one?

OpenStudy (anonymous):

i get zero

OpenStudy (anonymous):

let cos x = u then you have u^2, d/dx(u^2) = 2u * u', 2* cosx * - sinx = -2cosxsinx

OpenStudy (anonymous):

can you use the equations please. its difficult to read and understand how you wrote it

OpenStudy (bahrom7893):

2*Cos(x)*(-Sinx) - (-Sin(2x)*2)

OpenStudy (anonymous):

\[(\cos ^{2}x - \cos(x ^{2}))'\] For \[\cos ^{2}x\] Let u = cosx, Then \[\cos ^{2}x\] = \[u ^{2}\] Since it is a sum/difference of functions, we can take the derivatives seperately, so the derivative of \[u ^{2}\] is 2u * u', which is 2cosx * - sinx

OpenStudy (anonymous):

\[\cos x ^{2}\] = \[\sin x^{2} * (x^{2})'\] You use the chain rule again, you take the derivative of the outer function(cosx), then the inner function (x^2)

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

-sin* my bad